An operation manager in charge of a company's manufacturing keeps track of the number of LED television in a day. Compute for the following data that represents the number of the LED television manufactured for the past three (3) weeks: 20,18,19,25,20,21,20,25,30,29,28,29,25,25,27,26,22. a. Construct frequency distribution (frequency,cumulative frequency, class boundaries and midpoint). In determining the classes of the data you will use (Rules #1). b. Compute for the mean,median and mode of the data using grouped data. c. Construct a stem and leaf of the data.


1
Expert's answer
2022-04-04T17:27:54-0400

Class Midpoint(X) Frequency (F) Cumulative Frequency

18-21 19.5 6 6

22-25 23.5 5 11

26-29 27.5 5 16

30-33 31.5 1 17

Total 17

mean ="\frac{\sum fx}{\sum f}"

Mean ="\frac{19.5\cdot6+23.5\cdot 5+27.5\cdot5+31.5\cdot1}{17}" =23.74

Median =L+ "\frac{n/2-c}{f}\cdot h"

Median =17.5+"\frac{9.25-6}{6}\cdot3" =19.125

Mode =L+ "\frac{f1-f0}{2f1-fo-f2}\cdot h"

Mode=17.5+"\frac{17.5-0}{12-0-5} \cdot 3" =25



Stem and leaf

1 8 9

2 0 0 0 1 2 5 5 5 5 6 7 8 9 9

3 0